OpenStudy (anonymous):

x + 3y = -4 4x + 5y = 5 solve using substitution

OpenStudy (anonymous):

4x + 5y = 5 =>4(-3y-4) + 5y = 5 =>-12y-16+5y=5 =>7y=-21 =>y=-3 =>x=5